40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 255 of 336  Not logged in ELOG logo
ID Date Author Type Category Subject
  4132   Tue Jan 11 11:19:13 2011 josephbSummaryCDSStoring FE harddrives down Y arm

Lacking a better place, I've chosen the cabinet down the Y arm which had ethernet cables and various VME cards as a location to store some spare CDS computer equipment, such as harddrives.  I've added (or will add in 5 minutes) a label "FE COMPUTER HARD DRIVES" to this cabinet.

  4131   Tue Jan 11 01:05:29 2011 kiwamuUpdateLSCobtained Xarm PDH signal

[Jenne, Zach, Kiwamu]

 

 We made some efforts to lock the X arm cavity with the infrared beam.

We eventually obtained the PDH signal from a photo diode at AS port, we are still in the mid way of the lock.

The PDH signal now is going into c1lsc's ADC.

We have to make sure which digital channel corresponds to our PDH signal,

 

(what we did)

- split the LO signal, which comes from a Marconi, just before the EOM into two path.

     One is going to the EOM and the other is going to the AP table for the demodulation, The driving frequency we are using is 11MHz.

- put RF amplifiers to make the RF signal bigger. The raw signal was small, it was about ~-50 dBm in the spectrum analyzer.

   So we connected two ZLN amplifiers. Now the RF signal is at about 0 dBm

- connected the LO and RF signals to a mixer.  Additionally we put a 9.5-11.5 MHz bandpass filter at the LO path since there was some amounts of 29.5MHz due to the RF reflection at the EOM resonant box.

     After a low pass filter by SR560 the signal shows typical PDH behaviors.

- strung a BNC cable which connects the demodulated signal and c1lsc.

    In order to connect the signal to the current working ADC, we disconnected AS166_I from a whitening board and plug our cable on it.

- tried checking the digital signal but we somehow couldn't configure DAQ setting, So actually we couldn't make sure which channel corresponds to our signal.

 

  4130   Mon Jan 10 16:47:08 2011 josephb, alex, rolfUpdateCDSFixed c1lsc dolphin reflected memory

While Alex and Rolf were visiting, I pointed out that the Dolphin card was not sending any data, not even a time stamp, from the c1lsc machine.

After some poking around, we realized the IOP (input/output processor) was coming up before the Dolphin driver had even finished loading. 

We uncommented the line

#/etc/dolphin_wait

in the /diskless/root/etc/rc.local file on the frame builder.  This waits until the dolphin module is fully loaded, so it can hand off a correct pointer to the memory location that the Dolphin card reads and writes to.  Previously, the IOP had been receiving a bad pointer since the Dolphin driver had not finished loading.

So now the c1lsc machine can communicate with c1sus via Dolphin and from there the rest of the network via the traditional Ge Fanuc RFM.

  4129   Mon Jan 10 16:39:36 2011 josephb, alex, rolfUpdateCDSNew Time server for frame builder and 1PPS

Alex and Rolf came over today with a Tempus LX  GPS network timing server.  This has an IRIG-B output and a 1PPS output.  It can also be setup to act as an NTP server (although we did not set that up).

This was placed at waist height in the 1X7 rack.  We took the cable running to the presumably roof mounted antenna from the VME timing board and connected it to this new timing server.  We also moved the source of the 1PPS signal going to the master timer sequencer (big blue box in 1X7 with fibers going to all the front ends) to this new time server.  This system is currently working, although it took about 5 minutes to actually acquire a timing signal from the GPS satellites.  Alex says this system should be more stable, with no time jumps. 

I asked Rolf about the new timing system for the front ends, he had no idea when that hardware would be available to the 40m.

Currently, all the front ends and the frame builder agree on the time.  Front ends are running so the 1 PPS signal appears to be working as well.

  4128   Sun Jan 9 15:50:55 2011 ranaHowToPSLSetting the PMC gain

Untitled.png

I ramped the PMC gain slider to find where it oscillates. It starts going bad at ~13 dB, so the new default gain is 7 dB to give us some margin for alignment improvements, etc.


I also fixed the TIME field in our MEDM screens by adding the following text to the C1IFO_STATE.db file which runs on c1iscaux:
grecord(stringin, "C0:TIM-PACIFIC_STRING")
{
    field(DESC, "Current time and date")
    field(DTYP, "EPICS IOC VAR")
    field(SCAN, "1 second")
    field(INP, "C1:FEC-34_TIME_STRING")
}

grecord(stringin, "C0:IFO-TIME_PACIFIC")
{
    field(DESC, "Current time and date")
    field(DTYP, "EPICS IOC VAR")
    field(SCAN, "1 second")
    field(INP, "C1:FEC-34_TIME_STRING")
}

This gets the time info from the c1ioo processor via channel access and gives it these mroe reasonable names. The first record is for backwards compatibility. The second record is a better name and we should use it in the future for all new screens. I had to reboot c1iscaux several times to figure out the right syntax, but its OK now. You have to reopen stale screens to get the field to refresh.

This avoids the previous idea of changing all of the MEDM screens.

  4127   Sat Jan 8 23:18:03 2011 ranaUpdateComputersnetwork table

There's, as usual, a disconnect between the real martian host table and the one displayed on the Wiki. Basically, the point is: DO NOT TRUST THE WIKI, because people who update the actual host table only randomly update the wiki table.

The only thing worth trusting is the file

/var/named/chroot/var/named/martian.zone
on linux 1.

You will need to have 'root' or 'sudo' to get into that directory to read the file.

  4126   Sat Jan 8 21:12:12 2011 ranaUpdateCDSMegatron is back

 I started reverting Megatron into a standard Ubuntu workstation after Joe/Osamu's attempt to steal it for their real time mumbo jumbo.

First, I installed a hard drive that was sitting around on top of it. That whole area is still a mess; I'm not surprised that we have so many CDS problems in such a chaotic state. There's another drive sitting around there called 'RT Linux' which I didn't use yet.

Second, I removed the ethernet cables and installed a monitor/keyboard/mouse on it.

Then I popped in the Ubuntu 10.04 LTS DVD, wiped the existing CentOS install and started the standard graphical installation of Ubuntu.

megatron.jpg

Megatron's specs attached: 

Attachment 2: sysinfo.text
  4125   Fri Jan 7 15:17:37 2011 steveUpdateCamerasITMX video monitor has tower wide view

The Watec 902, 1/2" CCD camera got new Tamron lens: 1/2" 10-40mm F/1.4  manual iris. IR corrective lens. It is designed to have the same focal point in the IR as

in the visible light range. However, as the depth of field in the IR range is very narrow, focus adjustment should be done carefully in the IR.

Now you can see the sus tower that will make alignment easier.

 

  4124   Fri Jan 7 12:01:39 2011 OsamuUpdateComputersc1lsc running

I got a new adapter board for expansion chassis from CDS and exchanged the existing adapter board which was laid on the floor around ETMX to new one.

Then I connected the chassis to c1lcs,  c1lsc seems to be running now. I will return the old board to CDS since Rolf says he wants to return it to manufacture.

 I found an interface box from ADC to D-SUB37pin and a cable to connect them.

I needed to make cables to connect the interface box to existing LSC whitening filters that has a 37 pin female D-SUB connector on one end and a 40pin female flat connector on the other end. We should use shielded cables for them, but unfortunately CDS did not have right one. Temporarily I made one cable for 1-8ch using a ribbon twist cable like Joe did.

I found a saturation at ch5 of ADC0 on c1lsc. I did not check carefully but it seemed to come from the LSC whitening board. Input of ch5 of the whitening board was not terminated and had a huge output voltage, but also ch6 was not terminated and had no big output. I guess something wrong on the LSC whitening board. Needs to be checked. Anyway I unplugged the small ribbon cable between the whitening board and the next LSC AA filter board.

Finally I realized that fiber connection of RFM did not exist. What I saw was the fiber cable of Dolphin. We need a RFM PCIe interface board, and a long fiber cable between c1lsc and RFM hub.

  4123   Fri Jan 7 00:49:16 2011 JenneUpdateGreen LockingRecovered Xarm Green Lock, Preparation for Beat Note Measurement

[Kiwamu, Jenne]

We went this evening in search of a beat note signal between the Xarm transmitted green light and the PSL doubled green light. 

First, we removed our new ETMX camera (we left the mount so it should be easy to put back) from the other day.   We left the test masses exactly where they had been while flashing for IR, so even though we can no longer confirm, we expect that the IR beam axis hasn't changed.  We used the steering mirror on the end table to align the green beam to the cavity.  We turned on the loop to lock the end laser to the cavity, and achieved green lock of the arm.

Then we went to the PSL table to overlap the arm transmitted light with the PSL doubled light.  We made a few changes to the optics that take the arm transmitted light over to the PD.  We found that the arm transmitted light was very high, so we changed from having one steering mirror to having 3 (for table space / geometry reasons we needed 3, not just 2) in order to lower the beam axis.  We also found that the spot size of the arm transmitted beam was ~2x too small, so we changed the mode matching telescope from a 4x reducer to a 2x reducer by changing the 2nd lens from f=50mm to f=100mm (the first lens is f=200mm).  We made the arm transmitted beam and the PSL green beam overlap, but we saw no peak on the spectrum analyzer. 

We checked the temperature of the PSL and end lasers, and determined that we needed to adjust the set temp of the end laser.  However, we still didn't find any beat note.

We then tweaked the temperature of the doubling oven at the end station, to maximize the power transmitted, since Kiwamu said that that had worked in the past.  Alas, no success tonight.

We're stopping for the evening, with the success of reacquiring green lock of the Xarm.

  4122   Fri Jan 7 00:14:36 2011 kiwamuUpdateIOOEOM triple resonant box is working

 I checked the triple resonant box for the broadband EOM this afternoon, and found it was healthy.

So I installed it again together with a power combiner and succeeded in locking the MC.

Since the box has a non-50 Ohm input impedance at 29.5 MHz, so it maybe needed to adjust the phase of the LO for the demodulation of the PDH signal.

A good thing is that now we are able to impose the other sidebands (i.e. 11 MHz and 55MHz) via the power combiner.

  4121   Thu Jan 6 10:47:11 2011 KojiUpdateelogELOG fixed (re: restarted)

Fixed the 40m elog crashing with the threaded display.

This morning I found that Google bot crashed the elog again. I started the investigation and found the threaded mode is fine if we use the recent 10 entries.

I gradually copied the old entries to a temporary elog and found that a deleted elog entry on August 6 had a corrupted remnant in the elog file. This kept crashed the threaded mode.

Once this entry has been eliminated again, the threaded mode got functional.

I hope this eliminates those frequent elog crashing.

Quote:

Google bot  crashed the elog again. Then, I found that Google bot (and I) can crash elogd by trying to show the threaded view.
There looks similar issue reported to the elog forum, the author did not think this is a true bag.

Note: This happens only for the 40m elog. The other elogs (ATF/PSL/TCS/SUS/Cryo) are OK for the threaded view.

Quote:

Did it again. It seemed that Google bot came to the elog and tried to obtain "http://nodus.ligo.caltech.edu:8080/robots.txt". That was the last of the log.
Bot came from the AJW's homepage. Also Google FeedFecther came to the elog.

 

 

  4120   Thu Jan 6 00:06:01 2011 JenneUpdateIOOMagical absorbing PZT mirror

[Kiwamu, Jenne]

We have a measely 465mW going into the MC.  We lose a boatload of power on the PZT mirror that is part of the last zigzag for steering into the MC.  Just before this mirror, we measure 1.21W .  Just after this mirror, we measure ~475mW.  Then a teeny bit gets picked off for the PSL POS/ANG.  But we're losing a factor of 3 on this one mirror.  Need to fix!!!!!!!!!

  4119   Thu Jan 6 00:03:05 2011 KojiUpdateCamerasAligned the Xarm, no big deal

Nice, nice!

The power budget for the FPMI is here. The expected Intracavity power and the transmission are at  most ~8W and 100uW, respectively.

 

  4118   Wed Jan 5 23:45:33 2011 JenneUpdateCamerasAligned the Xarm, no big deal

Correct.  We can see the flashes clearly on our new ETM camera, but we see absolutely nothing on the ITM.

Unfortunately, the camera is in the path of the green beam, so we'll have to figure out a more permanent plan.  Right now the laser at the end is shuttered.

Measuring the power now....

Quote:

Pretty fancy work actually...I wouldn't have guessed that you need to look at the back of the ETM. Does this mean that we can't see the flashes from inside the arm cavity? If so, its a pretty remarkeable statement about the wide angle scatter of the new mirrors. Do we really have 1W going into the MC?

 

  4117   Wed Jan 5 23:37:12 2011 ranaUpdateCamerasAligned the Xarm, no big deal

Pretty fancy work actually...I wouldn't have guessed that you need to look at the back of the ETM. Does this mean that we can't see the flashes from inside the arm cavity? If so, its a pretty remarkeable statement about the wide angle scatter of the new mirrors. Do we really have 1W going into the MC?

  4116   Wed Jan 5 23:22:00 2011 JenneUpdateCamerasAligned the Xarm, no big deal

[Kiwamu, Jenne]

We successfully aligned the X arm.  No big deal.  Nothing to write in giant colorful letters about.  If we thought it was tricky, we'd be excited.  But since we're rockstar grad students, we can do this anytime, with one hand behind our back.

The details:

Earlier this evening, Kiwamu put a PD at the dark port.  After starting with the usual steering beams around and approximately centering them by finding the beam on the SUS tower, we saw that we could see the fringes on a 'scope hooked up to the dark port's new PD.  We could make the dip in the scope trace go away by misaligning the ETM, so we were confident that it was due to some kind of resonance in the arm.  We then fine-tuned our beam centering by moving the optics in either pitch or yaw until the fringe went away, wrote down the number, then moved the other direction until the fringe went away, and then put the optic back in the middle of those two numbers.  We did the ETM first, then the ITM (because the beam on the PD is sensitive to the ITM pointing, so we didn't want to have to move the ITM very far).  We saw that the cavity had a visibility of ~56% when we had finished with this method.

We then went to look for the flashes transmitted through the ETM.  We were not able to see them on a card, but when we looked with an IR viewer at the back face of the ETM, we could see the flashes. We stole a spare CCD camera found on the PSL table, and the camera power supply from the RefCav Refl camera, and set up a CCD camera with telephoto lens on the ETMX Trans table, looking directly at the back of the ETM.  We hooked the camera up to the regular ETMX camera cable, so we can see the flashes in the control room.  You can see them here:

DSC_2822.JPG

While the cavity was aligned, here were the slider positions:

KiwamuJenne_5Jan2011_Xarm_Aligned.png

  4115   Wed Jan 5 22:14:41 2011 ranaSummaryDAQFrameBuilder fails in a new way

Just a proof that the DAQ is working - ran DTT on nodus from 3 hours ago.

Attachment 1: Screen_shot_2011-01-05_at_10.13.21_PM.png
Screen_shot_2011-01-05_at_10.13.21_PM.png
  4114   Wed Jan 5 21:19:29 2011 ZachUpdateelogrestarted

 Restarted the elog with the script

  4113   Wed Jan 5 16:11:17 2011 kiwamuSummaryIOOtemporary PZT connection

PZTconnection.png

This is a connection diagram for the input PZTs (i.e. PZT1 and PZT2).

As drawn in the diagram, the signals don't go through the anti-imaging filter D000186 in the current configuration.

  4112   Wed Jan 5 16:00:11 2011 rana, alexSummaryDAQFrameBuilder fails in a new way

Email from Alex:

Turned out to be the lack of current year information in the IRIG-B signal
received by the Symmetricom GPS card in the frame builder machine caused
this. I have added a constant in daqdrc to bring the seconds forward:

controls@fb /opt/rtcds/caltech/c1/target/
fb $ grep symm daqdrc
#set symm_gps_offset=-1;
set symm_gps_offset=31536001;

Hopefully we will be upgrading to the newer timing system at the 40M this
year, so this will not happen again next year.


 

Doing an 'ls -lrt' in /frames/full/ now shows that the names are correct:

drwxr-xr-x 2 controls controls 249856 Dec 30 19:06 9777
drwxr-xr-x 2 controls controls 192512 Dec 31 16:00 9778
drwxr-xr-x 2 controls controls 151552 Jan  1 08:53 9463
drwxr-xr-x 2 controls controls 266240 Jan  2 12:39 9464
drwxr-xr-x 2 controls controls 258048 Jan  3 16:26 9465
drwxr-xr-x 2 controls controls 253952 Jan  4 20:13 9466
drwxr-xr-x 2 controls controls 151552 Jan  5 13:54 9467
drwxr-xr-x 2 controls controls  12288 Jan  5 15:57 9783

  4111   Wed Jan 5 10:45:51 2011 KojiUpdateelogrestarted

Google bot  crashed the elog again. Then, I found that Google bot (and I) can crash elogd by trying to show the threaded view.
There looks similar issue reported to the elog forum, the author did not think this is a true bag.

Note: This happens only for the 40m elog. The other elogs (ATF/PSL/TCS/SUS/Cryo) are OK for the threaded view.

Quote:

Did it again. It seemed that Google bot came to the elog and tried to obtain "http://nodus.ligo.caltech.edu:8080/robots.txt". That was the last of the log.
Bot came from the AJW's homepage. Also Google FeedFecther came to the elog.

 

  4110   Wed Jan 5 03:06:02 2011 kiwamuUpdateIOObeam spots on MC mirrors

I checked the spot positions on MC1 and MC3 by running Yuta's A2L script.

The amounts of the off-centering were good except for YAW of MC1.

 So we have to adjust the YAW alignment of the beam axis by steering the mirrors at the PSL table.

- - - (measured off-centering) - - - 

MC1_PIT  =  -0.711 mm

MC1_YAW =  1.62 mm

MC3_PIT  =   -0.0797 mm

MC3_YAW  =  - 0.223 mm

 

Quote:

We will check the spot positions more accurately by A2L technique.

 

  4109   Wed Jan 5 00:23:30 2011 ranaSummaryDAQFrameBuilder fails in a new way

Since Leo was trying to demo his LIGO Data Listener code, he noticed that there was and NDS2 issue. The NDS2 guy (JZ) noticed that the FrameBuilder had an issue.

We investigated. At 4PM on Dec 31, the GPS timestamp of the frame file names started to be recorded wrong. In fact, it started to give it a file name matching the correct time from 1 year in the past.

So that's our version of the Y2011 bug. Here's the 'ls' of /frames/full:

drwxr-xr-x 2 controls controls 252K Dec 26 03:59 9773
drwxr-xr-x 2 controls controls 260K Dec 27 07:46 9774
drwxr-xr-x 2 controls controls 256K Dec 28 11:33 9775
drwxr-xr-x 2 controls controls 252K Dec 29 15:19 9776
drwxr-xr-x 2 controls controls 244K Dec 30 19:06 9777
drwxr-xr-x 2 controls controls 188K Dec 31 16:00 9778
drwxr-xr-x 2 controls controls 148K Jan  1 08:53 9463
drwxr-xr-x 2 controls controls 260K Jan  2 12:39 9464
drwxr-xr-x 2 controls controls 252K Jan  3 16:26 9465
drwxr-xr-x 2 controls controls 248K Jan  4 20:13 9466
drwxr-xr-x 2 controls controls  36K Jan  5 00:22 9467
controls@fb /frames/full $

The culprit is the directory who's name starts out as 9463 whereas it should be 9779.

 

  4108   Tue Jan 4 21:21:57 2011 kiwamuUpdateIOOPZTs are connected to c1iscaux

I connected PZT1 and PZT2 to a slow front end machine c1iscaux.

Now we are able to align these PZTs from the control room via epics.

 

   Since we removed C1ASC that was controlling the voltage applied on the PZTs, we didn't have the controls for them for a long time.

So Rana and I decided to hook them up to an existing slow front end machine temporarily.

(probably the best solution is to connect them to C1LSC, which is fast enough to dither them.)

We actually found that c1iscaux is the proper machine, because it looked like it used to control the PZTs a long long time ago.

Moreover, c1iscaux still has DAC channels named like C1:LSC-PZT1_X, and so on.

 

  Below shows a screen shot of the medm screen for controlling the PZTs, invoked from a button on sitemap.adl ( pointed by a black arrow in the picture below)

The current default values are all zero at the right top sliders.

Screenshot20110104.png

  4107   Tue Jan 4 18:37:18 2011 JenneUpdateIOOMCWFS aligned

I undid Yuta's temporary setup, and put beam back on both WFS.  Since Koji had just aligned the Mode Cleaner, I centered the beam on the WFS using the WFS QPD screen, while watching the WFS Head screen, to make sure that the beam was actually hitting the QPD, and not off in lala land. 

Quote from Koji:

- We must check the MCWFS path alignment and configuration.

Quote from Jenne:

* Noticed that the MCWFS path is totally wrong.  Someone (Yuta?) wanted to use the MCWFS as a reference, but the steering mirror in front of WFS1 was switched out, and now no beam goes to WFS2 (it's blocked by part of the mount of the new mirror). I have not yet fixed this, since I wasn't using the WFS tonight, and had other things to get done.  We will need to fix this.

 

  4106   Tue Jan 4 15:12:33 2011 SureshUpdateIOOMC Trans Mon QPD gain decreased by 10

 Decreased the gain of MC-Trans-Mon QPD ckt

The resistors R1, R2, R3, R4 are now 49.9 kOhm. The previous elog on this subject 3882 has the ckt details. 

  4105   Tue Jan 4 13:57:48 2011 kiwamuUpdateIOOincident mirror changed from 45deg to 0 deg

I replaced the final steering mirror (Y1-1037-45-S) in the zig-zag path on the PSL table by a 0 deg mirror Y1-1037-0.

With a sensor card I confirmed the transmission reduced a lot after the replacement.

As we expected, the replacement of the mirror caused a mis-alignment of the incident beam axis to the MC, so I compensated it by touching the angle of the mirror a little bit.

After the alignment of the mirror, the MC is still resonating at TEM00.

We will check the spot positions more accurately by A2L technique.

Quote:

- Y1-45S has 4% of transmission. Definitively we like to use Y1-0 or anything else. There must be the replaced mirror. 

I think Suresh replaced it. So he must remember wher it is.

  4104   Tue Jan 4 11:06:32 2011 KojiUpdateIOOPower into Mode Cleaner increased

- Previously MC TRANS was 9000~10000 when the alignment was good. This means that the MC TRANS PD is saturated if the full power is given.
==> Transimpedance must be changed again.

- Y1-45S has 4% of transmission. Definitively we like to use Y1-0 or anything else. There must be the replaced mirror.
I think Suresh replaced it. So he must remember wher it is.

- We must confirm the beam pointing on the MC mirrors with A2L.

- We must check the MCWFS path alignment and configuration.

- We should take the picture of the new PSL setup in order to update the photo on wiki.

Quote:

What was the point:

I twiddled with several different things this evening to increase the power into the Mode Cleaner.  The goal was to have enough power to be able to see the arm cavity flashes on the CCD cameras, since it's going to be a total pain to lock the IFO if we can't see what the mode structure looks like.

Summed-up list of what I did:

* Found the MC nicely aligned.  Did not ever adjust the MC suspensions.

* Optimized MC Refl DC, using the old "DMM hooked up to DC out" method.

* Removed the temporary BS1-1064-33-1025-45S that was in the MC refl path, and replaced it with the old BS1-1064-IF-2037-C-45S that used to be there.  This undoes the temporary change from elog 3878.  Note however, that Yuta's elog 3892 says that the original mirror was a 1%, not 10% as the sticker indicates. The temporary mirror was in place to get enough light to MC Refl while the laser power was low, but now we don't want to fry the PD.

* Noticed that the MCWFS path is totally wrong.  Someone (Yuta?) wanted to use the MCWFS as a reference, but the steering mirror in front of WFS1 was switched out, and now no beam goes to WFS2 (it's blocked by part of the mount of the new mirror). I have not yet fixed this, since I wasn't using the WFS tonight, and had other things to get done.  We will need to fix this.

* Realigned the MC Refl path to optimize MC Refl again, with the new mirror.

* Replaced the last steering mirror on the PSL table before the beam goes into the chamber from a BS1-1064-33-1025-45S to a Y1-45S.  I would have liked a Y1-0deg mirror, since the angle is closer to 0 than 45, but I couldn't find one.  According to Mott's elog 2392 the CVI Y1-45S is pretty much equally good all the way down to 0deg, so I went with it.  This undoes the change of keeping the laser power in the chambers to a nice safe ~50mW max while we were at atmosphere.

* Put the HWP in front of the laser back to 267deg, from its temporary place of 240deg.  The rotation was to keep the laser power down while we were at atmosphere.  I put the HWP back to the place that Kevin had determined was best in his elog 3818.

* Tried to quickly align the Xarm by touching the BS, ITMX and ETMX.  I might be seeing IR flashes (I blocked the green beam on the ETMX table so I wouldn't be confused.  I unblocked it before finishing for the night) on the CCD for the Xarm, but that might also be wishful thinking.  There's definitely something lighting up / flashing in the ~center of ETMX on the camera, but I can't decide if it's scatter off of a part of the suspension tower, or if it's really the resonance. 

What happened in the end:

The MC Trans signal on the MC Lock screen went up by almost an order of magnitude (from ~3500 to ~32,000).  When the count was near ~20,000 I could barely see the spot on a card, so I'm not worried about the QPD.  I do wonder, however, if we are saturating the ADC. Suresh changed the transimpedance of the MC Trans QPD a while ago (Suresh's elog 3882), and maybe that was a bad idea? 

Xarm not yet locked. 

Can't really see flashes on the Test Mass cameras. 

 

  4103   Tue Jan 4 02:58:53 2011 JenneUpdateIOOPower into Mode Cleaner increased

What was the point:

I twiddled with several different things this evening to increase the power into the Mode Cleaner.  The goal was to have enough power to be able to see the arm cavity flashes on the CCD cameras, since it's going to be a total pain to lock the IFO if we can't see what the mode structure looks like.

Summed-up list of what I did:

* Found the MC nicely aligned.  Did not ever adjust the MC suspensions.

* Optimized MC Refl DC, using the old "DMM hooked up to DC out" method.

* Removed the temporary BS1-1064-33-1025-45S that was in the MC refl path, and replaced it with the old BS1-1064-IF-2037-C-45S that used to be there.  This undoes the temporary change from elog 3878.  Note however, that Yuta's elog 3892 says that the original mirror was a 1%, not 10% as the sticker indicates. The temporary mirror was in place to get enough light to MC Refl while the laser power was low, but now we don't want to fry the PD.

* Noticed that the MCWFS path is totally wrong.  Someone (Yuta?) wanted to use the MCWFS as a reference, but the steering mirror in front of WFS1 was switched out, and now no beam goes to WFS2 (it's blocked by part of the mount of the new mirror). I have not yet fixed this, since I wasn't using the WFS tonight, and had other things to get done.  We will need to fix this.

* Realigned the MC Refl path to optimize MC Refl again, with the new mirror.

* Replaced the last steering mirror on the PSL table before the beam goes into the chamber from a BS1-1064-33-1025-45S to a Y1-45S.  I would have liked a Y1-0deg mirror, since the angle is closer to 0 than 45, but I couldn't find one.  According to Mott's elog 2392 the CVI Y1-45S is pretty much equally good all the way down to 0deg, so I went with it.  This undoes the change of keeping the laser power in the chambers to a nice safe ~50mW max while we were at atmosphere.

* Put the HWP in front of the laser back to 267deg, from its temporary place of 240deg.  The rotation was to keep the laser power down while we were at atmosphere.  I put the HWP back to the place that Kevin had determined was best in his elog 3818.

* Tried to quickly align the Xarm by touching the BS, ITMX and ETMX.  I might be seeing IR flashes (I blocked the green beam on the ETMX table so I wouldn't be confused.  I unblocked it before finishing for the night) on the CCD for the Xarm, but that might also be wishful thinking.  There's definitely something lighting up / flashing in the ~center of ETMX on the camera, but I can't decide if it's scatter off of a part of the suspension tower, or if it's really the resonance.  Note to self:  Rana reminds me that the ITM should be misaligned while using BS to get beam on ETM, and then using ETM to get beam on ITM.  Only then should I have realigned the ITM.  I had the ITM aligned (just left where it had been) the whole time, so I was making my life way harder than it should have been.  I'll work on it again more today (Tuesday). 

What happened in the end:

The MC Trans signal on the MC Lock screen went up by almost an order of magnitude (from ~3500 to ~32,000).  When the count was near ~20,000 I could barely see the spot on a card, so I'm not worried about the QPD.  I do wonder, however, if we are saturating the ADC. Suresh changed the transimpedance of the MC Trans QPD a while ago (Suresh's elog 3882), and maybe that was a bad idea? 

Xarm not yet locked. 

Can't really see flashes on the Test Mass cameras. 

  4102   Mon Jan 3 10:32:27 2011 kiwamuSummaryelogfound hung - restarted

Found exactly the same error messages at the end of the log file.

Quote: #4098

This was the error today:

GET /40m/ HTTP/1.1
Host: nodus.ligo.caltech.edu:8080
User-Agent: Mozilla/5.0 (Macintosh; U; Intel Mac OS X 10.6; en-US; rv:1.9.2.3) Gecko/20100401 Firefox/3.6.3
Accept: text/html,application/xhtml+xml,application/xml;q=0.9,*/*;q=0.8
Accept-Language: en-us,en;q=0.5
Accept-Encoding: gzip,deflate
Accept-Charset: ISO-8859-1,utf-8;q=0.7,*;q=0.7
Keep-Alive: 115
Connection: keep-alive
Referer: http://www.ligo.caltech.edu/~ajw/40m_upgrade.html
Cookie: elmode=threaded; __utma=65601905.411937803.1291369887.1291369887.1291369887.1; __utmz=65601905.1291369887.1.1.utmcsr=(direct)|utmccn=(direct)|utmcmd=(none); SITESERVER=ID=4981c5fd42ae53c9c9e0980f2072be4f

 

  4101   Sat Jan 1 19:13:40 2011 ranaUpdateCDSc1pem now recording data

 I found that there was no PEM data nor any other data (no SUS or otherwise. No testpoints, no DAQ).

I went through the procedure that Jenne has detailed in the Wiki but it didn't work.

1) Firstly, the 'telnet fb 8088' step doesn't work. It says "Connected to fb.martian" but then just hangs. To replicate the effect of this step I tried ssh'ing to fb and doing a 'pkill daqd'. That works to restart the daqd process.

2) The wiki instructions had a problem. In the GUI step, it should say 'Save' after the Acquire bit has been set to 1. Even so, this works to get the .ini file right and the DTT can see the correct channel list, but none of the channels are available. There are just 'Unable to obtain measurement data'.

3) I tried running 'startc1pem', but no luck. I also tried rebooting c1sus from the command line. That worked so far as to come back up with all the right processes running, but still no data. The actual /frames directory shows that there are frames, but we just can't see the data. I also tried to get data usind the DTT-NDS2 method, but still no luck. (*** ITMX and ITMY both came back with all their filters off; worth checking if their BURTs are working correctly.)

Using DataViewer, however, I AM able to see the data (although the channel name is RED). In fact, I am able to see the trend data ever since I changed the Acquire bit to 1. Plot attached as evidence. Why does DTT not work anymore???

Attachment 1: Untitled.png
Untitled.png
  4100   Sat Jan 1 16:36:16 2011 ranaConfigurationSUSrampdown.pl was not running

The crontab for op340m which runs various IFO maintenance activities has been set to the wrong path for the watchdog rampdown script since the CDS changeover.

This is a dangerous situation. With the watchdog thresholds set as high as 1000, the magnets can be broken if the new CDS has some mental problems. This kind of thing happened before at LHO (which is why Stan invented the watchdogs). Please try to make sure that the watchdog thresholds are not set that way - its our only defense against bad CDS.

I've now corrected the crontab. The watchdog thresholds are being stepped down every 30 minutes as before.

However, out test points are gone again, so who knows how things are behaving.

  4099   Fri Dec 31 12:41:02 2010 ranaUpdateGeneralIR flashes

I came in today to check up on the StripTool and burn the Ubuntu LTS (new CDS OS) DVD. I was pretty excited to see the PRM flashes on Mon1.

I waggled the PRM/BS alignments and got a good contrast MICH and then bright flashes in the PRM that totally overload Mon1's CCD.

Now I can see flashes of some IR junk in the X Arm; its way off on the left edge of the mirror, but there's a beam.

For the short term, we can hook up the IO PZTs to some old EPICS channels (like one of the AUX guys in the LSC area), but eventually it has to get hooked up to the new ASC or ASS. We have to bug Joe to see where this shows up in his master diagram.

**Note: if you get lost sometimes when doing the alignments, remember that you can use time_machine_conlog.

Ex:

rossa:general>./time_machine_conlog 2010/12/31,11:00:00 PDT C1:SUS-ITMX_PIT_COMM
Issuing the conlog command:
/cvs/cds/caltech/conlog/bin/conlog +epics -interp at 2010/12/31,11:00:00 PDT "C1:SUS-ITMX_PIT_COMM"


Conlog Replied:

LIGO controls: values at 2010 12/31 11:00:00 pst
 __Epics_Channel_Name______   __value_____
 C1:SUS-ITMX_PIT_COMM         0.406100



Restoring now...

Restoration Successful

Attached is the last 8 days of Vacuum Pressure trend which includes the pumpdown.

Attachment 1: Untitled.png
Untitled.png
  4098   Wed Dec 29 18:53:11 2010 ranaSummaryelogfound hung - restarted

This was the error today:

GET /40m/ HTTP/1.1
Host: nodus.ligo.caltech.edu:8080
User-Agent: Mozilla/5.0 (Macintosh; U; Intel Mac OS X 10.6; en-US; rv:1.9.2.3) Gecko/20100401 Firefox/3.6.3
Accept: text/html,application/xhtml+xml,application/xml;q=0.9,*/*;q=0.8
Accept-Language: en-us,en;q=0.5
Accept-Encoding: gzip,deflate
Accept-Charset: ISO-8859-1,utf-8;q=0.7,*;q=0.7
Keep-Alive: 115
Connection: keep-alive
Referer: http://www.ligo.caltech.edu/~ajw/40m_upgrade.html
Cookie: elmode=threaded; __utma=65601905.411937803.1291369887.1291369887.1291369887.1; __utmz=65601905.1291369887.1.1.utmcsr=(direct)|utmccn=(direct)|utmcmd=(none); SITESERVER=ID=4981c5fd42ae53c9c9e0980f2072be4f

  4097   Fri Dec 24 09:01:33 2010 josephbUpdateCDSBorrowed ADC

Osamu has borrowed an ADC card from the LSC IO chassis (which currently has a flaky generation 2 Host interface board).  He has used it to get his temporary Dell test stand running daqd successfully as of yesterday.

This is mostly a note to myself so I remember this in the new year, assuming Osamu hasn't replaced the evidence by January 7th.

  4096   Thu Dec 23 22:13:50 2010 KojiUpdateGeneralPZT HV turned on

The four IOO PZTs have been turned on in order to confirm the alignment of the IFO.

Once they are turned on, the spots (ITMX/ITMY/PRM/SRM) on the REFL CCD have been easily found.

When the X-arm was aligned to the green beam, it is easily locked to TEM00. Also some LG modes were visible.
i.e. There is some room to improve the mode matching.
The transmitted green at the PSL table is a bit too high and clipped by the first mirror on the table.

No IR flashes were found in either arms.

------------------

The below are the range and the set values of the strain gauge readback for the PZTs.
When the closed loop buttons are activated the PZTs are fixed at those values, if no one touches the set point dials.

            Min   Max   SetP | Display on the module
PZT1 Yaw    2.20  9.95  6.08 | Broken
PZT1 Pitch -0.011 8.89  4.40 | 1.58

PZT2 Yaw    0.737 9.94  5.37 | 2.17
PZT2 Pitch  0.010 9.42  4.71 | 1.89

  4095   Thu Dec 23 18:51:37 2010 ranaUpdateGreen Lockinginstalled doubling oven base at PSL table

This is not such a bad base design, but remember that we have to get a couple of parts with the purple anodize to see if there is a difference between black and purple in terms of the 532 nm reflectivity.

  4094   Thu Dec 23 13:30:09 2010 steveConfigurationVACslow pumpdown complete

The pump down continued this morning by the removal of the butterfly valve. Two roughing pumps were used to reach 500 mTorr

The Maglev monitoring MEDM screen "Rotating" indicator is not working. It is on all times. Please look at Maglev controller monitor for real information.

Pump down is completed.           

Configuration: vacuum normal after 86 days at atm               CC1 = 1e-5 Torr                  

                      IFO is hungry for light (and maybe some goulash with a little paprikash too)

Attachment 1: slow2dpd.jpg
slow2dpd.jpg
  4093   Thu Dec 23 12:19:43 2010 kiwamuUpdateGreen Lockinginstalled doubling oven base at PSL table

I gave a christmas present to a doubling oven who has been sitting on the PSL table.

The below is a picture of the present I gave. It's a base plate for the doubling oven, made from a block of aluminum, and black-anodized.

The size and the shape are nicely tailored for the combination of the Newfocus kinematic mount and the Covesion oven.

The design had been done by using Solid Works 2010

 

DSC_2820_ss.jpg

 

Here is a picture before he got the present.

DSC_2814_ss.jpg

 

Now he looks pretty happy.

DSC_2815_ss.jpg

  4092   Thu Dec 23 08:54:32 2010 SureshUpdateelogthe delhi univ syndrome

Sorry folks!  I couldnt get to the elog and didnt know that the elog was crashing every time I tried to access it. 

But have found other means to access it and the elog is safe for now!

  4091   Thu Dec 23 03:15:11 2010 KojiUpdateelogelogd is getting killed by Suresh

ELOG has crashed and I restarted it.

Actually the filtering is not effective so far as elog is not using apache but has its own web server inside.
So this just block the access to port 30889 (=SVN, Dokuwiki, etc).

Quote:

After another elog crash, I've blacklisted the domain that Suresh is using by editing the apache httpd.conf. Let's see what happens now.

 

  4090   Wed Dec 22 21:38:47 2010 kiwamuConfigurationVACRe:slow pumpdown at 12 hours

I stopped the pumping at 9:30 pm. Now we are at 29 torr.

Quote: from  #4089

How to stop pumping:

1,  close RV1 manual valve with torque wheel

2, close V3

3, turn off roughing pumps RP1 & RP3

4, disconnect metal hose connection to butterfly valve

 

  4089   Wed Dec 22 17:24:50 2010 steveConfigurationVACslow pumpdown at 12 hours

We have reached 200 Torr at 12 hours of slow pumping speed. Kiwamu stopped the pumping for 11 hrs last night .and I restarted it this morning.

Now RV1 is fully open with butterfly valve in place  and the second roughing  pump RP3 was just turned on.

 

How to stop pumping:

1,  close RV1 manual valve with torque wheel

2, close V3

3, turn off roughing pumps RP1 & RP3

4, disconnect metal hose connection to butterfly valve

Attachment 1: slowpd12h.jpg
slowpd12h.jpg
  4088   Wed Dec 22 15:47:22 2010 KojiUpdateVACTank closing procedure

[Jenne Kiwamu Joe Osamu Steve Koji]

Yesterday (21st), we closed the BSC and the four testmass chambers.

We splitted the team into three.

- Wiping team (Jenne/Kiwamu)
- Suspension team (Osamu/Koji)
- Door closing team (Joe/Steve)

While the wiping team was working, the suspension team prepared the next suspension.
Once the mirror was wiped, the suspension team restored that suspension at the position of the markings on the table.

After the wipe of the first mirror, the wiping team got experienced and they were the fastest among the teams.
So the wiping team worked as the second suspension team when necessary.

All of us became the closing team after the last suspension has been restored.

We checked that we still have the Michelson fringes and the oplev pointings.

In total, the work was very efficient such that we only took four hours for the entire process.

We left the suspension marking on the optical tables because they are quite useful.

  4087   Wed Dec 22 15:30:47 2010 OsamuUpdateASCETMY oplev fixed

According to c1scy.mdl, OL signals should be connected to adc_0_24 to adc_0_27 but they were connected to adc_0_16 to adc_0_19 which are assigned to QPD signals.

Actually cable connections were messed up. One ribbon cable was connected from QPD driver and ADC ports assigned for OL, and another ribbon cable was connected from the board combining the signals of oplev and QPD to ADC port assigned for QPD.

Now ETMY oplev is working well and aligned to center.

  4086   Wed Dec 22 11:24:23 2010 haixingUpdateSUSmeasurement of imbalance in quadrant maglev protope

Yesterday, a sequence of force and gain measurement was made to determine the imbalance in the

quadrant, magnetic-levitation prototype. This was the reason why it failed to achieve a stable levitation.

The configuration is shown schematically by the figure below:

 

config_2x2.png

 

Specifically, the following measurements have been made:

(1) DC force measurement among four pairs of magnets at fixed distance with current of the coils on and off

From this measurement,  the DC force between pair of magnets is determined and is around 1.6 N at with a

separation of 1 cm. This measurement also lets us know the gain from voltage to force near the working point.

 

The force between pair "2" is about 13% stronger than other pairs which are nearly identical. The force by the

coil is around 0.017 N per Volt (levitation of 5 g per 3 Volt); therefore, we need around 12 volt DC compensation

of pair "2" in order to counterbalance such an imbalance.  Given the resistence of the coil equal to 26 Om, this

requires almost 500 mA DC compensation. Koji suggested that we need a high-current buffer, instead of what

has been used now.

 

(2) DC force measurement among four pairs of magnets (with current of the coils off) as a function of distance

From this measurement, we can determine the stiffness of the system. In this case, the stiffness or the

effective spring constant is negative, and we need to compensate it by using a feedback control. This is

one of the most important parameters for designing the feedback control. The data is still in processing.

 

(3) Gain measurement of the OSEM from the displacement to voltage.

This measurement is a little bit tricky due to the difficulty to determine the displacement of the flag.

After several measurements, it gave approximately 2 V/cm.

 

Plan for the next few days:

From the those measurements, all the parameters for the plant and sensor that need to determine the

feedback control are known. They should be plugged into the simulink model and to see whether the

old design is appropriate or not. Concerning the experimental part, we will first try to levitate the configuration

with 2 pairs of magnets, instead of 4 pairs, as the first step, which is easier to control but still interesting.

 

 

 

 

 

 

  4085   Wed Dec 22 06:50:19 2010 ranaUpdateelogelogd is getting killed by Suresh

After another elog crash, I've blacklisted the domain that Suresh is using by editing the apache httpd.conf. Let's see what happens now.

  4084   Tue Dec 21 16:34:42 2010 kiwamuSummaryVACAll the test masses have been wiped

 [Jenne and Kiwamu]

 We wiped all the test masses with isopropyl alcohol.

They became apparently much cleaner.

(how to)

 At first we prepared the following stuff:

  * syringe

  * isopropyl alcohol 

  * lens papers

  * cleaned scissors

  Then we cut the lens papers into the half by the scissors such that the long side can remain.

This is because that the SOSs have relatively narrow spaces at their optic surfaces for putting a piece of paper. 

   We did vertical and horizontal wiping using the lens paper and appropriate amount of isopropyl alcohol.

Each wiping (vertical and horizontal) requires two or three times trials to appropriately remove dusts.

Amount of isopropyl:

   * vertical 15 [ul]

   * horizontal 10 [ul]

In addition to this, we also used the ionizer gun for blowing big dusts and fiber away from the surface.

 

 

(surface inspection)

   Before wiping them, all the test masses had small dusts uniformly distributed on the HR surfaces.

Especially ETMX was quite dirty, many small spots (dusts) were found when we shined the surface with the fiber illuminator.

ETMY was not so bad, only a couple of small dusts were at around the center.  ITMX/Y had several dusts, they were not as dirty as ETMX, but not cleaner than ETMY.

   After we wiped them,  we confirmed no obvious dusts were around the centers of the optics. They looked pretty good !

 

 

  4083   Tue Dec 21 16:20:09 2010 steveConfigurationVACpumpdown has started

Slow pumpdown of rebuilt IFO-mark4 started at 3pm today.

We are at 680Torr at this minute with 1.5 Torr/min speed.

ELOG V3.1.3-